Calculate the number of atoms in 3 moles of water of hydrogen and oxygen seperately and total number of atoms
1
Expert's answer
2020-07-07T14:37:09-0400

Solution:

Water (H2O) is made up of one oxygen atom and 2 hydrogen atoms.

So,

1 mol H2O = 2 mol H atoms = 1 mol O atoms


Moles of O atoms = Moles of H2O = 3 mol

Moles of H atoms = 2 × Moles of H2O = 2 × 3 mol = 6 mol


One mole of any substance contains 6.02×1023 units of that substance.

Thus:

The number of oxygen atoms = (3 mol) × (6.02×1023) = 18.06×1023 O atoms

The number of oxygen atoms = (6 mol) × (6.02×1023) = 36.12×1023 H atoms

The total number of atoms = No. of O atoms + No. of H atoms = (18.06 + 36.12)×1023 = 54.18×1023


Answer:

No. of O atoms is 18.06×1023

No. of H atoms is 36.12×1023

The total number of atoms is 5.42×1024